# RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation

### v1

@@ 4.1 Benchmarks
We use HumanEval (+) Chen et al. (2021) and MBPP(+) Austin et al. (2021); Liu et al. (2023) to assess the fundamental coding ability of code model. Meanwhile, we use LiveCodeBench Jain et al. (2024) and BigCodeBench-hard Zhuo et al. (2024) to measure the advanced coding ability of code model. We restrict data usage to the range from 240701 to 250201 of the LiveCodeBench to maintain its contamination-free nature.
Models Dataset HumanEval MBPP LCB BCB-hard Average
Base Plus Base Plus 2407-2502 Inst Comp
Proprietary Models
[…]
Claude-3.5-Sonnet-20240620 / 87.2 81.7 89.4 74.3 32.0 25.7 33.1 60.5
GPT-4o-20240806 / 92.7 87.2 87.6 72.2 30.5 25.0 36.5 61.7
Open-source Models 7B+ Scale
[…]
DeepSeek-V3 / 91.5 86.6 87.6 73.0 36.3 27.7 39.9 63.2
Llama3.3-70B-Instruct / 84.1 80.5 87.6 73.5 29.1 28.4 28.4 58.8
Open-source Models $\sim$7B Scale
[…]
DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 12.8 10.1 15.5 46.4
MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 13.4 13.5 12.8 48.0
OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 7.3 13.5 16.9 47.1
WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 12.9 12.8 16.9 46.5
RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 13.6 18.2 10.1 46.6
RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 14.1 18.9 14.2 47.7+1.1
RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 14.3 19.6 14.2 48.5+1.9
RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 14.2 20.3 15.5 49.0+2.4
[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9 / / 16.2      /
Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 18.1 20.3 20.3 55.2
RefineCoder-QW-7B   Iter0 20K 85.4 78.0 79.4 65.1 18.2 18.2 20.9 52.2
RefineCoder-QW-7B   Iter1 20K 86.0 79.3 84.7 71.4 17.4 16.2 18.9 53.4+1.2
RefineCoder-QW-7B   Iter2 20K 86.6 80.5 84.7 72.0 19.2 18.2 19.6 54.4+2.2
RefineCoder-QW-7B   Iter3 20K 87.2 81.1 85.2 73.0 19.1 19.6 20.9 55.2+3.0
Table 1: Main Results. LCB/BCB denotes LiveCodeBench/BigCodeBench. Inst/Comp denote Instruct/Complete. [Uncaptioned image] and [Uncaptioned image] denotes the two base models, and the rest are the instruction models. 2BT denotes 2B tokens.
@@ 4.2 Baselines
$\circ$ Proprietary Models: Gemini-1.5-Pro-002 Google (2024), Claude-3.5-Sonnet-20240620 Anthropic (2024), and GPT-4o-20240806 OpenAI (2024a).
$\circ$ Open-source Models 7B+ Scale: Qwen2.5-Coder-32B-Instruct Hui et al. (2024), DeepSeek-V3 DeepSeek-AI (2024), and Llama3.3-70B-Instruct Grattafiori et al. (2024).
$\circ$ Open-source Models $\sim$7B Scale: DSCoder-6.7B Series Guo et al. (2024), MagiCoder-S-DS-6.7B Wei et al. (2024b), OpenCodeInterpreter-DS-6.7B Zheng et al. (2024a), WaveCoder-Ultra-6.7B Yu et al. (2024), and Qwen2.5-Coder-7B Series Hui et al. (2024).
@@ 4.3 Initial SFT Dataset $D_{0}$
The existing datasets vary in quality and carry the risk of data leakage Wang et al. (2024b); Yu et al. (2024). To ensure the accuracy and reliability of the experimental results, we constructed a high-quality SFT dataset by combining the Evol-Instruct and SelfCodeAlign Wei et al. (2024a). Concretely, we first prompt GPT-4o to generate code-related concepts and corresponding programming instructions from Python code snippets. These code snippets have been preprocessed to ensure high quality and contamination-free. Then, we prompt GPT-4o again to iteratively evolve these instructions using Evol-Instruct strategy. Finally, we prompt GPT-4o to generate code responses for the instructions. Through this pipeline, we obtained an 80K high-quality and diverse python instruction dataset. See Appendix C for detailed data construction prompts.
Dataset Size HumanEval (+) MBPP (+) LCB BCB-hard
Magic-OSS-Instruct Wei et al. (2024b)  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K 75K 4.45 9.40 2.63 4.84
Magic-Evol-Instruct Wei et al. (2024b)  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K 110K 43.20 19.40 2.91 4.46
Evol-CodeAlpaca-v1 Luo et al. (2024)  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1 110K 47.04 19.46 2.91 4.46
Code-FeedBack Zheng et al. (2024a)  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback 68K 30.50 17.67 3.16 4.78
Ours 80K 4.97 7.00 1.54 5.43
Table 2: The Test Leakage Indicator (TLI) quantifies data leakage by measuring the average maximum n-gram overlap between dataset samples and benchmark samples. The larger the TLI value, the more severe the data leakage. Values in red indicate severe data leakage.
@@ 4.4 Implement Details
Iterative Training We employ DSCoder-6.7B-Base and Qwen2.5-Coder-7B-Base as the base pre-trained models. To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B (Iter0). Furthermore, since Qwen2.5-Coder-Base exhibits strong code understanding and generation capabilities, we accelerate our experimental iterations by randomly sampling 20K examples from the same dataset for fine-tune it, yielding RefineCoder-QW-7B (Iter0). The two RefineCoder models undergo three iterative training using their respective datasets under our ACR framework. In the ACR pipeline, the model self-samples 5 code responses with a temperature of 0.7. We use Qwen2.5-32B-Instruct Hui et al. (2024) as the Judge and Critic. We set the number of training epochs to 2, the global batch size to 64, and the learning rate to 5e-6, employing the AdamW optimizer with a cosine learning rate decay strategy. All our training is performed using 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al. (2024b).
Evaluation We use the Pass@1 metric to evaluate the performance of model on the benchmark. For baseline results, we prioritize those from the benchmark leaderboard  [^5] https://evalplus.github.io/leaderboard.html  [^6] https://livecodebench.github.io/leaderboard.html  [^7] https://huggingface.co/spaces/bigcode/bigcodebench-leaderboard or the original papers. If unavailable, we evaluate them locally using the same settings as the RefineCoder series.
@@ 4.5 Experimental Results
As shown in Table 1, the RefineCoder series achieves impressive performance gains on various code generation benchmarks. While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.4 and 3.0, respectively, on all benchmarks after three iterations of refinement. Furthermore, it is worth noting that the RefineCoder series is fine-tuned with a very limited amount of data (80K or 20K). Despite this, the RefineCoder models (Iter3) still outperform or match baselines of the same size, which are fine-tuned with more extensive data.
Concretely, data refinement improves the fundamental programming abilities of the RefineCoder series, as demonstrated by continuous performance gains over iterations on HumanEval, HumanEval+, and MBPP+. Notably, RefineCoder-QW-7B shows a remarkable 7.9 point increase (Iter0 -> Iter3) in Pass@1 on MBPP+, highlighting the efficacy of this approach. In the more challenging benchmarks, LiveCodeBench and BigCodeBench-hard, the performance of the RefineCoder also improves after iterations. Compared to the baseline OpenCodeInterpreter, RefineCoder-DS-6.7B (Iter2) outperforms it by 7 points on LiveCodeBench, illustrating the strong competitive programming abilities of the RefineCoder series. These results demonstrate that the model can iteratively enhance code generation capabilities by self-generated code and critique with the help of our ACR method.
@@ 5.1 Data Leakage Analysis
In Table 1, we observe that while the RefineCoder (Iter3) surpasses or matches the baselines on average pass@1, it still lags behind the state-of-the-art baseline of the same size on HumanEval (+) and MBPP+. To better explore this performance gap, we investigate potential data leakage in the datasets used by the baselines.
Concretely, we collect the open-source datasets used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (the datasets for the other baselines are not public) and evaluate data leakage using the TLI metric proposed by Wang et al. (2024b). The results are shown in Table 2, Magic-OSS-Instruct and our dataset are entirely free from data leakage. In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage with respect to HumanEval (+) and MBPP (+). These datasets are used by the three aforementioned baselines, making it challenging for RefineCoder to surpass these baselines. We also present a intuitive scatter plot of similarity scores in the appendix E for a more detailed analysis.
@@
RefineCoder-DS-6.7B Iter3 65.3 60.1-5.2 6.3
RefineCoder-QW-7B Iter3 70.4 69.6-0.8 5.8
Table 3: Best and Worst denote the performance of code with the highest and lowest scores. Error rate denotes the proportion of cases where the best code fails the test cases, but the worst code passes.
@@ 5.2 The effectiveness of Composite Scoring System
The performance improvement through iterative training preliminarily validates the effectiveness of the scoring system. In this section, we further verify it by enabling the model sampling 10 code responses for each programming question in benchmarks and scoring them. As shown in Table 3, the performance of the highest-scoring code far exceeds that of the lowest-scoring code, while the scoring system maintains a low error rate. This clearly validates its effectiveness.
5.3 The effectiveness of Selective Critique Strategy
To validate the effectiveness of the selective critique strategy in ACR, we conduct ablation experiments. As shown in Figure 3, for each iteration, we construct two variants of the RefineCoder without the 3.1 module in Figure 2, and another without the 3.2 module. During the iteration process, ablating either module leads to a performance drop in both RefineCoder-QW-7B and RefineCoder-DS-6.7B, indicating the effectiveness of our strategy. Furthermore, removing the second-round critique data has a greater impact on performance. We believe this action prevents the model from reflecting on self-generated low-quality codes and external critiques, thereby hindering performance improvement.
Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on all benchmarks.
@@ 5.4 Multilingual Code Generation
Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
@@
RefineCoder-QW-7B   Iter0 60.2 73.4 68.4 49.4 78.0 78.3 68.7
RefineCoder-QW-7B   Iter1 62.7 74.1 65.2 51.3 76.1 77.0 68.6
RefineCoder-QW-7B   Iter2 63.8 74.1 69.6 48.7 78.0 77.6 69.4
RefineCoder-QW-7B   Iter3 60.2 75.3 73.4 53.2 79.2 77.6 70.6
@@
As shown in Table 4, we also evaluate the out-of-distribution (OOD) code generation performance on multilingual benchmark MultiPL-E Cassano et al. (2022), despite fine-tuning on a Python-only dataset. After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 1.9 points, respectively. In particular, RefineCoder-QW-7B (Iter3) achieves best results among half of the programming languages. This demonstrates that the effectiveness of our ACR method generalizes well to the OOD code generation domain.
@@ 5.5 Evaluation with External Feedback
Iterative ACR not only improves the one-off code generation performance of RefineCoder but also endows it with the ability to correct errors based on feedback. Following  Zheng et al. (2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor. 2) Human Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking. Model then refines the code based on these suggestions. The results of RefineCoder-QW-7B on HumanEval (+) and MBPP (+) are shown in Figure 4. It can be observed that external feedback improves performance on all benchmarks, with human feedback yielding a more significant enhancement. Surprisingly, with the help of human-like improvement suggestions, RefineCoder-QW-7B (Iter1 and Iter3) outperforms GPT-4o on three benchmarks, demonstrating its ability to understand feedback and refine code.
Models HumanEval MBPP
Base Plus Base Plus
Qwen2.5-Coder-7B-Base 61.6 53.0 76.9 62.9
     +DPO 59.8 52.4 44.2 32.6
     +ORPO 85.4 79.9 84.4 70.4
RefineCoder-QW-7B Iter0 85.4 78.0 79.4 65.1
     +DPO 61.0 50.0 60.6 42.3
     +ORPO 86.6 79.3 81.7 69.3
Table 5: Performance comparison of different preference optimization strategies in reinforcement learning.
5.6 Preference Optimization
During the iterative process, a substantial amount of preference data is generated. For each programming instruction, we select the top-1 and bottom-1 ranked responses to form preference pairs, thereby exploring the potential of preference learning. As shown in Table 5, we use DPO Rafailov et al. (2023) and ORPO Hong et al. (2024) to optimize the base and SFT models. We observe the opposite results due to the differences in preference strategies. DPO leads to a significant performance decline due to the intrinsic risk of reducing the likelihood of correct code during training Feng et al. (2024); Tao et al. (2024); Pal et al. (2024). In contrast, ORPO builds upon the SFT loss to maximize the likelihood of correct code, resulting in improved performance. The results are consistent with those presented in the CodeLutra Tao et al. (2024), reflecting the indispensability of SFT loss and the critical impact of preference strategy design on performance.

### v2

@@ 4.1 Benchmarks
We use HumanEval(+) Chen et al. (2021), MBPP(+) Austin et al. (2021); Liu et al. (2023) and BigCodeBench Zhuo et al. (2024) to assess the fundamental coding ability of code model. Meanwhile, we use LiveCodeBench Jain et al. (2024) and BigCodeBench-hard to measure the advanced coding ability of code model.
@@ 4.2 Baselines
$\circ$ Based on DSCoder-6.7B-Base: DSCoder-6.7B-Instruct Guo et al. (2024), MagiCoder-S-DS-6.7B Wei et al. (2024b), OpenCodeInterpreter-DS-6.7B Zheng et al. (2024a), WaveCoder-Ultra-6.7B Yu et al. (2024).
$\circ$ Based on Qwen2.5-Coder-7B-Base: Qwen2.5-Coder-7B-Instruct Hui et al. (2024).
@@ 4.3 Initial SFT Dataset $D_{0}$
The existing datasets vary in quality and carry the risk of data leakage Wang et al. (2024b); Yu et al. (2024). To ensure the accuracy and reliability of the experimental results, we construct a high-quality SFT dataset by combining Evol-Instruct and SelfCodeAlign Wei et al. (2024a). Concretely, we first prompt GPT-4o to generate code-related concepts and corresponding programming instructions from Python code snippets. These code snippets have been preprocessed to ensure high quality and contamination-free. Then, we prompt GPT-4o again to iteratively evolve these instructions using the Evol-Instruct strategy. Finally, we prompt GPT-4o to generate code responses for the instructions. Through this pipeline, we obtain an 80K high-quality and diverse python instruction dataset. See Appendix B for detailed data construction prompts.
Models Data Size HumanEval MBPP BigCodeBench Average
Base Plus Base Plus Inst Comp
[Uncaptioned image] DSCoder-6.7B-Base / 47.6 39.6 72.0 58.7 / 41.8 /
DSCoder-6.7B-Instruct 2BT 74.4 71.3 74.9 65.6 35.5 43.8 60.9
MagiCoder-S-DS-6.7B 185K 76.8 71.3 79.4 69.0 36.2 47.6 63.4
OpenCodeInterpreter-DS-6.7B 178K 77.4 72.0 76.5 66.4 37.1 44.6 62.3
WaveCoder-Ultra-6.7B 130K 75.0 69.5 74.9 63.5 33.9 43.7 60.1
RefineCoder-DS-6.7B   Iter0 80K 73.8 67.7 77.5 65.1 37.0 46.6 61.3
RefineCoder-DS-6.7B   Iter1 80K 74.4 68.9 77.0 66.4 40.6 48.0 62.6↑1.3
RefineCoder-DS-6.7B   Iter2 80K 74.4 70.3 79.6 66.9 40.6 47.9 63.3↑2.0
RefineCoder-DS-6.7B   Iter3 80K 75.0 70.7 80.2 67.2 41.1 49.6 64.0↑2.7
[Uncaptioned image]Qwen2.5-Coder-7B-Base / 61.6 53.0 76.9 62.9    /    /    /
Qwen2.5-Coder-7B-Instruct Millions 88.4 84.1 83.5 71.7 40.4 48.8 69.5
RefineCoder-QW-7B   Iter0 80K 84.9 78.1 79.4 65.8 40.0 47.4 65.9
RefineCoder-QW-7B   Iter1 80K 86.0 79.3 84.1 70.8 41.6 49.0 68.5↑2.6
RefineCoder-QW-7B   Iter2 80K 86.6 82.5 83.4 71.7 42.3 50.9 69.6↑3.7
RefineCoder-QW-7B   Iter3 80K 87.2 83.3 85.6 72.9 42.0 51.1 70.4↑4.5
Table 2: Results on HumanEval, MBPP and BigCodeBench. Values in red indicate severe data leakage.
Dataset Size HumanEval (+) MBPP (+) LCB BCB
Magic-OSS-Instruct Wei et al. (2024b) 75K 4.45 9.40 2.63 4.56
Magic-Evol-Instruct Wei et al. (2024b) 110K 43.20 19.40 2.91 4.10
Evol-CodeAlpaca-v1 Luo et al. (2024) 110K 47.04 19.46 2.91 4.69
Code-FeedBack Zheng et al. (2024a) 68K 30.50 17.67 3.16 4.49
Ours 80K 4.97 7.00 1.54 5.00
Table 3: The Test Leakage Indicator (TLI) quantifies data leakage by measuring the average maximum n-gram overlap between dataset samples and benchmark samples. The larger the TLI value, the more severe the data leakage.
@@ 4.4 Implement Details
Iterative Training We employ DSCoder-6.7B-Base and Qwen2.5-Coder-7B-Base as the base pre-trained models. To obtain an initial code model $M_{0}$, we fine-tune DSCoder-6.7B-Base and Qwen2.5-Coder-Base on the 80K initial SFT dataset, resulting in RefineCoder-DS-6.7B and RefineCoder-QW-7B. The two RefineCoder models undergo three iterative training under our ACR framework. In the ACR pipeline, the model self-samples 7 code responses with a temperature of 0.7. We use Qwen2.5-32B-Instruct Hui et al. (2024) as the Judge and Critic. We set the number of training epochs to 2, the global batch size to 64, and the learning rate to 5e-6, employing the AdamW optimizer with a cosine learning rate decay strategy. All our training uses 16 A100-80G GPUs, utilizing the LLaMA-Factory Zheng et al. (2024b).
Evaluation We use the Pass@1 metric to evaluate the performance of model on benchmarks. We prioritize baseline results from the benchmark leaderboards or the original papers. If unavailable, we evaluate them locally using the same settings as the RefineCoder series.
@@ 4.5 Experimental Results
As shown in Table 1, the RefineCoder series achieves impressive performance gains on two challenging code benchmarks. While the initial RefineCoder-DS-6.7B and RefineCoder-QW-7B exhibit moderate performance, the two models achieve an average Pass@1 improvement of 2.7 and 3.1, respectively, after three refinement iterations. Notably, the performance improved by 2.0 and 2.3 points from iter0 to iter1, highlighting that incorporating self-generated code and external critique yields better results than directly distilling code responses from the teacher model. Compared to baselines of the same size, the RefineCoder-DS-6.7B and RefineCoder-QW-7B (Iter3) outperform them by 3.6 and 2.3 points, respectively, using only 80K data. This demonstrates the superiority of our method in improving advanced coding ability.
Similarity, our method also improves the basic coding ability, as shown in Table 2. After three iterations, our model achieves average performance improvements of 2.7 and 4.5 points, respectively, on HumanEval(+), MBPP(+) and BigCodeBench. This demonstrates that the model is capable of identifying and addressing deficiencies by learning from both its self-generated code and external feedback. Compared with all baselines, our models achieve the best average Pass@1 performance. Nevertheless, we find that the performance of RefineCoders on the Humaneval(+) are inferior to the state-of-the-art baseline. This discrepancy prompts us to conduct a data leakage analysis of the datasets in the next section.
@@ 5.1 Data Leakage Analysis
In Table 2, we observe that while the RefineCoder (Iter3) surpasses the baselines on average pass@1, it still lags behind the state-of-the-art baseline on HumanEval(+) and MBPP+. To better explore this performance gap, we investigate potential data leakage in the datasets used by the baselines.
Concretely, we analyze the four open-source datasets  [^1] https://huggingface.co/datasets/ise-uiuc/Magicoder-OSS-Instruct-75K  [^2] https://huggingface.co/datasets/ise-uiuc/Magicoder-Evol-Instruct-110K  [^3] https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1  [^4] https://huggingface.co/datasets/m-a-p/Code-Feedback with the help of the TLI metric proposed by Wang et al. (2024b). These datasets are used by MagiCoder-S-DS-6.7B, OpenCodeInterpreter-DS-6.7B, and WaveCoder-Ultra-6.7B (The datasets for DSCoder-6.7B-Instruct and Qwen2.5-Coder-7B-Instruct are not publicly available). The results are shown in Table 3, Magic-OSS-Instruct and our dataset are entirely free from data leakage. In contrast, Magic-Evol-Instruct, Evol-CodeAlpaca-v1, and Code-FeedBack demonstrate severe data leakage concerning HumanEval (+) and MBPP (+). These three datasets are used by three baselines, making it challenging for RefineCoder to surpass them. We also present an intuitive scatter plot of similarity scores in the appendix D for a more detailed analysis.
@@
RefineCoder-DS-6.7B Iter3 70.3 60.1-10.2 5.8
RefineCoder-QW-7B Iter3 73.4 71.1-2.3 5.3
Table 4: Best and Worst denote the performance of code with the highest and lowest scores. Error rate denotes the proportion of cases where the best code fails the test cases but the worst code passes.
@@ 5.2 The effectiveness of Composite Scoring System
The performance improvement through iterative training preliminarily validates the effectiveness of our scoring system. In this section, we further verify it by enabling the model to sample 10 code responses for each programming question in benchmarks and scoring them. As shown in Table 4, the performance of the highest-scoring code far exceeds that of the lowest-scoring code, while the scoring system maintains a low error rate. This clearly validates its effectiveness.
Figure 3: Ablation Study. The y-axis denotes the average pass@1 value on LiveCodeBench and BigCodeBench-hard.
Figure 4: Evaluation with feedback using RefineCoder-QW-7B.
5.3 The effectiveness of Selective Critiquing Strategy
We conduct ablation experiments to validate the effectiveness of the selective critiquing strategy in ACR. As shown in Figure 3, for each iteration, we construct two variants of RefineCoder (without 3.1 module of Figure 2 or without 3.2 module). During the iteration process, ablating either module leads to a performance drop in both RefineCoder-QW-7B and RefineCoder-DS-6.7B, indicating the effectiveness of our strategy. Furthermore, removing the second-turn critique data has a greater impact on performance. We believe this action prevents the model from reflecting on self-generated low-quality code and external critique, thereby hindering performance improvement.
@@ 5.4 Multilingual Code Generation
RefineCoder-QW-7B   Iter0 60.7 72.4 68.0 49.9 77.6 78.1 67.8
RefineCoder-QW-7B   Iter1 62.1 74.6 65.0 50.8 76.8 77.4 67.8
RefineCoder-QW-7B   Iter2 63.7 74.1 69.6 49.7 78.3 77.7 68.9
RefineCoder-QW-7B   Iter3 60.9 75.3 73.8 52.2 79.3 77.9 69.9
@@
As shown in Table 5, we also evaluate the out-of-distribution (OOD) code generation performance on the multilingual benchmark MultiPL-E Cassano et al. (2022) despite fine-tuning on a Python-only dataset. After one and three iterations, RefineCoder-DS-6.7B and RefineCoder-QW-7B achieve optimal average performance, surpassing the initial model by 1.0 and 2.1 points, respectively. In particular, RefineCoder-QW-7B (Iter3) achieves the best results among half of the programming languages. This demonstrates that the effectiveness of our ACR method generalizes well to the OOD code generation domain.
@@ 5.5 Evaluation with External Feedback
Iterative ACR not only improves the one-off code generation performance of RefineCoder but also endows it with the ability to correct errors based on feedback. Following  Zheng et al. (2024a), we design two types of external feedback to evaluate this ability of RefineCoder: 1) Execution Feedback: Model refines its self-generated erroneous code based on execution results from the executor; 2) Human-like Feedback: GPT-4o first analyzes the programming question, initial error code, and execution results to generate improvement suggestions that mimic human thinking. The model then refines the code based on these suggestions. The results of RefineCoder-QW-7B on HumanEval(+) and MBPP(+) are shown in Figure 4. It can be observed that external feedback improves performance on all benchmarks, with human-like feedback yielding a more significant enhancement. Surprisingly, with the help of human-like improvement suggestions, RefineCoder-QW-7B (Iter1 and Iter3) outperforms GPT-4o on three benchmarks, demonstrating its ability to understand feedback and refine code.